OpenStudy (anonymous):

Multiply. Express your answer in lowest terms. -3/-8 * (-2/9) =?

Nnesha (nnesha):

\[\huge\rm \frac{ -3 }{ -8 } \times \frac{ -2 }{ 9 }\] like this ??

OpenStudy (anonymous):

No, The problem is like this:

Nnesha (nnesha):

ohh okay you typed -8 and -3 so that's why :) \[\huge\rm \frac{ -3 }{ 8 } \times \frac{ -2 }{ 9 }\] like this ?? now looks good

Nnesha (nnesha):

 divide 3 and 9 3 times 1 = 3 3 times 3 = 9 so cancel out that

OpenStudy (anonymous):

The 2/9 is negative like 3/8, Sorry for confusion.

Nnesha (nnesha):

nah its okay :) -:)

OpenStudy (anonymous):

So its 8?

Nnesha (nnesha):

nope first you can reduce fraction 3 goes into 9 how many times ?

OpenStudy (anonymous):

3 goes into 9, 3 times.

Nnesha (nnesha):

yep right so reduce fraction